How many grams of methamphetamine (C10H15N) would be needed to prepare 100 ml of a solution of 0.10 M methamphetamine?
1
Expert's answer
2018-04-26T07:42:09-0400
Answer:
Cm = n/V, where n is amount of substance, mole, V is volume of solution, litre, Cm is molar concentration, mole/litre
n = Cm*V
n = 0.1*0.1 = 0.01 mole
m = M*n, where M is molar mass (149 g/mole)
m = 149*0.01 = 1.49 g